Calculate the effect of an increase in pressure of
Answer #1

Benzene molar mass = 78 g/mol has been taken during calculations.

2ene Various tables Boiling point Tbf= 80,tc= 353.25 K Lipuid 0181 a Triple pont ft = 5, se = 278-65K critical point Tc= 562-

레8-8645 7076.25 pa = 53.1 Tow The ansoe helpr ou